2x-3y__=4
__-3y+z=2
x____-z= -5
How do you solve??? Please Hurry('X

Answer by lynnlo(4176)

if this is 3 different problems:
2x-3y=4=x3y/2+2
-3y+z=2=y=2/3-2/3
x-2=-5=x=-3